1. A golf ball is hit with an initial velocity of 50.0 m/s at an angle of 34 degrees above the horizontal. If the golf ball lands on the same horizontal level, determine the time it takes the ball to stay in the air.

2. A ball is thrown horizontally, with a speed of 10 m/s, from the top of a 4.9 meter tall hill. How far from the point on the ground directly below the launch point does the ball strike the ground?
Answer #1

time of flight

T = 2 u sin x / g

T = 2* 50* sin 34 / 9.8

T = 5.706 s

========

Considering motion along vertical

h = 0.5 gt^2

4.9 = 0.5 *9.8 t^2

t = 1 s

Distance travelled in t seconds

D = 10* 1 = 10 m

========

Comment in case any doubt, will reply for sure.. Goodluck